# Investigating the Event-Shape Methods in Search for the Chiral Magnetic Effect in Relativistic Heavy Ion Collisions

### v1

@@ Appendix A Simulation Plots
Figure A.11: Toy model ESS results. Toy Model I with the accepted $K_{S}$ fraction of the default 2% (left panel) and of 10% (middle panel), and Toy Model II with mass continuum (right panel). The $K_{S}$, $\rho^{0}$, and primordial pion inputs are from parameterizations to experimental data of 30–40% centrality Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ GeV. The numbers of events are $5\times 10^{6}$ for each panel. The $\Delta\gamma$ is plotted as a function of $v_{2,{\rm single}}$ in events binned in $\hat{q}_{2,{\rm pair}}^{2}\{2\}$ (Eqs. 5,9). POIs are from acceptance $|\eta|<1$ and $20.2<p_{\rm T}<$ 2 GeV/$c$, and the event selection variable $\hat{q}_{2,{\rm pair}}^{2}\{2\}$ is computed from the same POIs. The model’s known impact parameter direction $\psi=0$ is taken as the EP in calculating $\Delta\gamma$ (Eqs. 2,3) and $v_{2,{\rm single}}$ (Eq. 10).
@@
Figure A.12: Toy model ESE results. Toy Model I with the accepted $K_{S}$ fraction of the default 2% (left panel, $10^{10}$ events) and of 10% (middle panel, $5\times 10^{8}$ events), and Toy Model II with mass continuum (right panel, $10^{7}$ events). The $K_{S}$, $\rho^{0}$, and primordial pion inputs are from parameterizations to experimental data of 30–40% centrality Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ GeV. The $\Delta\gamma$ is plotted as a function of $v_{2}$ in events binned in $\hat{q}_{2}^{2}\{2\}$ (Eqs. 5,6). POIs are from acceptance $20.3<|\eta|<$ 2, and the event selection variable $\hat{q}_{2}^{2}\{2\}$ is computed from particles in $|\eta|<0.3$, both with $20.2<p_{\rm T}<$ 2 GeV/$c$. The model’s known impact parameter direction $\psi=0$ is taken as the EP in calculating $\Delta\gamma$ (Eqs. 2,3) and $\left\langle v_{2}\right\rangle$ (Eq. 10).

### v3

@@ Appendix A Simulation Plots
Figure A.12: Toy model ESS results. Toy Model I with the accepted $K_{S}$ fraction of the default 2% (left panel) and of 10% (middle panel), and Toy Model II with mass continuum (right panel). The $K_{S}$, $\rho^{0}$, and primordial pion inputs are from parameterizations to experimental data of 30–40% centrality Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ GeV. The numbers of events are $2\times 10^{8}$ for each panel. The $\Delta\gamma$ is plotted as a function of $v_{2,{\rm single}}$ in events binned in $\hat{q}_{2,{\rm pair}}^{2}\{2\}$ (Eqs. 5,9). POIs are from acceptance $|\eta|<1$ and $20.2<p_{\rm T}<$ 2 GeV/$c$, and the event selection variable $\hat{q}_{2,{\rm pair}}^{2}\{2\}$ is computed from the same POIs. The model’s known impact parameter direction $\psi=0$ is taken as the EP in calculating $\Delta\gamma$ (Eqs. 2,3) and $v_{2,{\rm single}}$ (Eq. 10). The red line is a first-order polynomial fit in the range of $0.20<v_{2,{\rm single}}<$ 0.2.
@@
Figure A.13: Toy model ESE results. Toy Model I with the accepted $K_{S}$ fraction of the default 2% (left panel, $2.8\times 10^{9}$ events) and of 10% (middle panel, $2.0\times 10^{9}$ events), and Toy Model II with mass continuum (right panel, $2.2\times 10^{8}$ events). The $\Delta\gamma$ is plotted as a function of $\left\langle v_{2}\right\rangle$ in events binned in $\hat{q}_{2}^{2}\{2\}$ (Eqs. 5,6). POIs are from acceptance $10.3<|\eta|<$ 1, and the event selection variable $\hat{q}_{2}^{2}\{2\}$ is computed from particles in $|\eta|<0.3$, both with $20.2<p_{\rm T}<$ 2 GeV/$c$. The model’s known impact parameter direction $\psi=0$ is taken as the EP in calculating $\Delta\gamma$ (Eqs. 2,3) and $\left\langle v_{2}\right\rangle$ (Eq. 10). Other information is as same as those in Fig. A.14. The red line is a first-order polynomial fit to all data points.
Figure A.14: Toy model ESE results (forward/backward rapidity $\hat{q}_{2}^{2}$). Toy Model I with the accepted $K_{S}$ fraction of 20% (left panel, $2.2\times 10^{8}$ events) and of 30% (middle panel, $2.2\times 10^{8}$ events), and Toy Model II with mass continuum (right panel, $10^{8}$ events). The event selection variable $\hat{q}_{2}^{2}\{2\}$ is computed from particles in $43<|\eta|<$ 4. Other information is as same as those in Fig. A.14.